OpenStudy (anonymous):

Simplify completely 12 times x to the third power minus 4 times x to the 2nd power plus 8 times x all over negative 2 times x

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

\(\Large\frac{12x^3-4x^2+8x}{-2x}\) Okay, for this one, you want to do the same thing - eliminate the denominator. Do you remember how to do that?

OpenStudy (anonymous):

I have 3 more after this thank u for helping me while everyone else just watchs and never helps

OpenStudy (anonymous):

no i do not get this

OpenStudy (anonymous):

i just need help with 3 more questions

OpenStudy (studygurl14):

I'll walk you through it one more time. First Step: factor out the the denominator from the numerator, in this case, -2x   So... \(\Large\frac{12x^3-4x^2+8x}{-2x}\rightarrow\frac{-2x(-6x^2+2x-4)}{-2x}\) What's the next step?